OpenStudy (anonymous):

Determine whether each sequence is arithmetic or geometric. 12, 8, 4, 0

OpenStudy (whpalmer4):

Arithmetic sequences have a common difference between successive terms. 12-8 = 8-4= 4-0= If those all give the same value, then it is arithmetic. Geometric sequences have a common ratio between successive terms. 12/8= 8/4= 4/0= If those all give the same value, then it is geometric.